Factoring:
a)x^10-x^8-2x^4-1
b)12x^2+5x-12y^2+12y-10xy-3
c)20x^4-9x^3-13x^2+7x-6
d)4(x^2+15x+50)(x^2+18x+72)-3x^2
Help me with those special types of factoring
You don't need to help me all.just which question you know .Thanks

Answer by ikleyn(52824)

.
a)  x^10-x^8-2x^4-1 = x^10 - (x^8 +2x^4 + 1) = x^10 - (x^4+1)^2 = 


    = (x^5 - (x^4+1))*(x^5 + (x^4+1)) = (x^5 - x^4 - 1)*(x^5 + x^4 + 1).